OpenStudy (anonymous):

-8x+8y=0 the answer was (0,0) and (7,7). How did they get this answer?

OpenStudy (anonymous):

that is a trivial function, -8x+8y=0 implies x = y so, the solution set are all the coordinates having the abscissas and ordinates the same.

OpenStudy (anonymous):

Sorry kimcarlos, I'm really doing my best to understand your response. I really want to follow but can you make it the reply simpler please?

OpenStudy (anonymous):

-8x+8y=0 -8x = -8y then, x=y so, if x=0 then y=0 implies (0,0) if x=7 then y=7 implies (7,7) hence you have coordinates (0,0) and (7,7)

OpenStudy (anonymous):

Thank you I see it now kimcarlos. What if I chose another number would it work the same? Like (5,5)

jimthompson5910 (jim_thompson5910):

like kimcarlos said "-8x+8y=0 implies x = y" so this means that as long as the x and y coordinate are the same, it's going to be a solution

jimthompson5910 (jim_thompson5910):

(5,5) would work because x = 5 and y = 5 so the x and y coordinates are the same as well

OpenStudy (anonymous):

thanks @jim_thompson5910 for the support

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

Thank you Jim. I've been struggling with the problems that end in 0. For example I had 6x-3y=0 but the answer wasn't completely different and I've been trying to understand the differences in the problems equaling 0.

OpenStudy (anonymous):

I meant to say the answer was completely different. I've been at this math since yesterday and it's beating me up

jimthompson5910 (jim_thompson5910):

I would solve that for y 6x-3y=0 6x-3y-6x=0-6x -3y = -6x -3y/(-3) = -6x/(-3) y = 2x so 6x-3y=0 is the same as y = 2x

jimthompson5910 (jim_thompson5910):

now plug in any value you want for x so let's say we do x = 5 y = 2x y = 2*5 y = 10 So when x = 5, y = 10 Therefore, (5,10) is a solution

OpenStudy (anonymous):

Okay Jim I'm going to practice this with other substitutions and see if I can be successful.

jimthompson5910 (jim_thompson5910):

alright sounds good

OpenStudy (anonymous):

I'm utilizing Hawkes Learning and the tutor is somewhat hard to follow on this software.

jimthompson5910 (jim_thompson5910):

is that a DVD set or something?

OpenStudy (anonymous):

It is a DVD program that is linked to college math via the internet. So, all the homework, quizzes and exams are generated and taken on this site. The site grades it as well.

jimthompson5910 (jim_thompson5910):

thought so, is it effective at all? I've never taken a DVD program before so I'm curious

OpenStudy (anonymous):

The problem I have is that I practice and practice my math and I do it for 5 -6hrs a day. Just practicing. I'm not afraid of it but I just try to practice all the rules which is way to many and I get really confused at times. But, it is a weakness that I'm enduring to overcome.

OpenStudy (anonymous):

It's by Pearson's Education a format of Higher Learning many professors use it for most core courses.

OpenStudy (anonymous):

It does all the grading for the professors. It has video's and mp3 downloads and study guides too. It is very useful but also very expensive too. My Biology class book with the access code DVD was 196.00 also another course that is done online.

jimthompson5910 (jim_thompson5910):

well just practice alone won't do it you need to be able to translate everything into terms that you understand and can connect to...just drilling it in won't be enough

jimthompson5910 (jim_thompson5910):

that's pretty expensive though for a DVD course...hopefully you have a teacher to ask questions as well?

OpenStudy (anonymous):

I try to practice and take notes as I go along. I'm not doing well in math thus far and I'm determined to do better and to be successful. You may think this is funny but I really like math but math doesn't like me. So, I keep trying to get a little better by practicing so that I can pass my quizzes and homework assignments. The website that we use says that we have to pass all the math homework with 75% in ordered to be certified as "Mastered" in any lessons. When you pass the homework it gives you a certificate for each lesson showing that you mastered that lesson.

jimthompson5910 (jim_thompson5910):

I see, so there's a somewhat steep curve to pass but then again, 75% isn't that bad and you'll definitely get better at it for sure, just keep practicing

OpenStudy (anonymous):

I shall. Okay Jim I have a new problem I would like to give it a shot with you if you don't mind?

OpenStudy (anonymous):

Please let me know where I have an error. -3x+4y=0 -3x(0)+4y=0 4y=3x How am I doing so far Jim?

jimthompson5910 (jim_thompson5910):

first solve for y -3x+4y=0 4y=3x y = 3/4x now you can plug in various values of x to get corresponding values of y

jimthompson5910 (jim_thompson5910):

for instance, if x = 4, then y = 3/4x y = (3/4)(4) y = 12/4 y = 3 so when x = 4, y = 3 we have this solution: (4,3)

OpenStudy (anonymous):

Ok I did get 3/4 as the answer. Next I plugged in the number 5 and I got the answer -15/4

jimthompson5910 (jim_thompson5910):

when x = 5, y is... y = 3/4x y = (3/4)(5) y = 15/4 so when x = 5, y is y =15/4

OpenStudy (anonymous):

Okay, I think I see I now have to plug in the number 4 which is the denominator.

jimthompson5910 (jim_thompson5910):

yes or any multiple of 4

jimthompson5910 (jim_thompson5910):

that will guarantee you get a whole number for y

OpenStudy (anonymous):

Okay I'm getting it and it feels good to see that I'm catching on. I have to do two hw assignments by 11:59 tonight and I've been at this since 6pm and haven't moved from my desk since.

jimthompson5910 (jim_thompson5910):

well that's good that it's making sense, even if it's taking a while But as long as it's making sense, that's all that matters really

OpenStudy (anonymous):

Yes, I'm trying to work it out now and to see if I can get through it with no mistakes.

jimthompson5910 (jim_thompson5910):

you're doing pretty well in my book

OpenStudy (anonymous):

@Charlotte0725 could you gice me that hawkes software...i want to try it

OpenStudy (anonymous):

give

jimthompson5910 (jim_thompson5910):

well Charlotte0725 said above "My Biology class book with the access code DVD was 196.00 also another course that is done online." so I'm guessing around $196, but idk

OpenStudy (anonymous):

jeezz so expensive

OpenStudy (anonymous):

kimcarlos I would love to give it to you but you have to have an access code to use it and you have to choose the professor who's teaching the course once you are signed on.